OpenStudy (anonymous):

What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6? please anyone help me

rishavraj (rishavraj):

if two lines are parallel ...it means their slopes are equal...so get the slope and apply \[\frac{ y_2 - y_1 }{ x_2 - x_1 } = m\]

OpenStudy (anonymous):

Is the slope 3/2?

rishavraj (rishavraj):

yup :))

OpenStudy (anonymous):

so know what do I do? sorry for being so ignorant but this is kinda hard for me

rishavraj (rishavraj):

now just use the equation \[\frac{ y_2 - y_1 }{ x_2 - x_1 } = m \] plug values

OpenStudy (anonymous):

But usnt that just to get the slope?

rishavraj (rishavraj):

nah u need to plug hmmmm like y_1 = 4 and y_2 = 2 so it would be \[\frac{ y_2 - 2 }{ x_2 - 4 } = \frac{ 3 }{ 2 }\]

rishavraj (rishavraj):

get equation in y_2 and x_2 and then replace x_2 by x and y_2 by y

OpenStudy (anonymous):

I got 3/2x-2

rishavraj (rishavraj):

nope @braandooll first of all cross multiply

OpenStudy (anonymous):

ohh so its Y=3/2x-4 huh?

rishavraj (rishavraj):

the equation is \[\frac{ y - 2 }{ x - 4 } = \frac{ 3 }{ 2 } \] so on cross multiplying \[2y - 4 = 3x - 12\] now proceed

OpenStudy (anonymous):

Thats why, i got.... y=3/2x-4 again

rishavraj (rishavraj):